# MultiGran-STGCNFog: Towards Accurate and High-Throughput Inference for Multi-Granular Spatiotemporal Traffic Forecasting

### v1

@@ Abstract
Accurate traffic forecasting and swift inference provision are essential for intelligent transportation systems. However, the present Graph Convolutional Network (GCN)-based approaches cannot extract and fuse multi-granular spatiotemporal features across various spatial and temporal scales sufficiently, proven to yield less accurate forecasts. Besides, additional feature extraction branches introduced in prior studies critically increased model complexity and extended inference time, making it challenging to provide fast inference for traffic forecasting. In this paper, we propose MultiGran-STGCNFog, an efficient fog distributed inference system with a novel traffic forecasting model that employs multi-granular spatiotemporal feature fusion on generated dynamic traffic graphs to fully capture interdependent traffic dynamics. The proposed scheduling algorithm GA-DPHDS, optimizing layer execution order and layer-device scheduling scheme simultaneously, contributes to considerable inference throughput improvement by leveraging heterogeneous fog devices in a pipelined manner. Extensive experiments on real-world datasets demonstrate the superiority of the proposed method over selected baselines.
1 Introduction
@@
With the advancement of Internet of Things (IoT) technology, making cities smarter through IoT has become a prominent research trend. The large volume of data collected by IoT devices has shifted attention toward data analysis algorithms and corresponding servings to enhance transportation system intelligence Zantalis et al. (2019). Traditional centralized transportation systems are limited by high latency and weak scalability. Against this backdrop, fog computing, which processes data on nearer nodes close to the data source, has emerged as a more efficient and scalable solution.
Having IoT sensors set up on roads and a large volume of traffic data gathered, such data needs to be analyzed efficiently to predict future traffic conditions. Accurate predictions benefit services like transportation optimization, navigation planning, and congestion prevention Li et al. (2022). Consequently, traffic forecasting has become an essential task, focusing on predicting road network patterns using historical observations Liu et al. (2023).
While GCN-based methods dominate spatiotemporal forecasting, transformer-based models have recently shown advantages in capturing long-term dependencies Jiang et al. (2023); Liu et al. (2023). However, GCNs still remain irreplaceably unmatched in handling spatial dependencies, and with carefully designed architectures, they are capable of achieving better performance than transformers Dan et al. (2024); Zhang et al. (2024). Considering their advanced technological maturity and broad applicability in spatiotemporal traffic-related tasks, we focuse on GCN-based approaches.
@@
Existing GCN and transformer models primarily design complex Spatial-Temporal (S-T) blocks but lack a structured approach to extract and integrate hierarchical, multi-granular spatiotemporal features across spatial and temporal dimensions. Although a few works have explored multi-granular spatial Wang et al. (2021); Guo et al. (2021) or temporal features Wang and Jing (2022); Huo et al. (2023); Guo et al. (2019) independently, seamlessly integrating both is critical for accurate forecasting and remains unexplored.
@@
Previous studies manipulating multi-scale traffic data introduced multiple feature extraction branches to capture features of different granularities. This significantly increased model complexity, leading to more extended inference time, which is challenging for Graph Neural Network (GNN) serving where efficiency and low latency are critical. No solutions were proposed in these earlier works to address this issue.
@@
Besides, designing accurate traffic forecasting models with swift inference still encounters several key challenges:
@@
1) Traffic data exhibits highly dynamic patterns, making it difficult to model spatiotemporal correlations.
@@
2) In GNN serving systems, efficiently leveraging available resources is essential for speeding up where communication latency becomes a major hurdle Zeng et al. (2022).
@@
3) Optimizing distributed parallel inference schemes is increasingly complex with the exponentially-growing search space and layer input-output dependencies.
@@
4) Mainstream distributed parallel frameworks, such as GPipe Huang et al. (2019), assume even partitioning of stages, ignoring hardware heterogeneity. Real-word device clusters are often heterogeneous, and model layers cannot be fully in even distribution to ensure equal processing times.
@@
To fill these gaps, we propose MultiGran-STGCNFog, an efficient GNN inference system that extracts and fuses multi-granular spatiotemporal features across various spatial and temporal scales. The distributed pipeline-parallel architecture across fog nodes enables MultiGran-STGCNFog to tackle the extended inference latency. These designs support swift traffic forecasting while enhancing forecasting performance.
@@
  • A novel GCN, titled Multi-Granularity Spatiotemporal Graph Convolution Network (MultiGran-STGCN), is developed. It employs dynamic graph techniques and a multi-granular spatiotemporal feature fusion mechanism through Laplacian-driven hierarchical graph clustering and multi-scale temporal modeling.
@@
  • To mitigate the prolonged inference latency induced by additional feature extraction branches in multi-scale modeling, we construct MultiGran-STGCNFog to accelerate inference, leveraging the resources of heterogeneous fog devices. It enables faster inference by pipeline parallelism, which allows immediate data transmission among fog nodes in proximity to avoid remote transfers.
@@
  • We develop a robust heterogeneous cross-device scheduling algorithm, GA-DPHDS, which jointly optimizes the layer execution order and the layer-device scheduling strategy, overcoming the search space explosion challenge. Our experiments demonstrate the necessity of layer execution order optimization, which is overlooked in previous studies with pipeline parallelism.
@@
  • The proposed framework is tested with physical computing nodes and evaluated using three real-world traffic datasets. Experimental results demonstrate its superior performance compared to selected baselines.
2 Related Work
2.1 GCNs: Spatiotemporal Traffic Forecasting
@@
When combined with temporal modeling techniques like RNNs, GCN-based models are particularly effective at capturing spatiotemporal dependencies Seo et al. (2018); Cui et al. (2020). DCRNN Li et al. (2017) modeled traffic as a diffusion process on directed graphs, using bidirectional diffusion convolution and RNNs to outperform baselines. Gated-STGCN Yu et al. (2017) leveraged graph convolutions for spatial and gated 1D convolutions for temporal modeling, improving efficiency by removing recurrent structures. ASTGCN Guo et al. (2019) employed attention mechanisms to capture spatial and temporal dependencies, while HGCN Guo et al. (2021) modeled traffic across micro- and macro-level networks. GWNET Wu et al. (2019) combined dilated causal and graph convolutions for hidden dependency capture, achieving state-of-the-art results, and OGCRNN Guo et al. (2020) integrated GCN with recurrent units for enhanced feature extraction.
2.2 Distributed and Parallel GNNs
@@
The growing computational demands of deep learning and the inability to deliver fast serving have driven a shift from centralized to distributed computing, leveraging parallelism in GNNs. Data parallelism partitions input data across nodes, using graph parallelism to minimize inter-node communication or mini-batch parallelism with node sampling. Model parallelism distributes the model itself through operator parallelism (simultaneous operations), pipeline parallelism (processing stages sequentially), or ANN parallelism (spreading layers across devices) Besta and Hoefler (2024).
@@
GPipe Huang et al. (2019) applied pipeline parallelism by dividing models into stages across accelerators for micro-batch processing, while PipeDream Narayanan et al. (2019) enhanced this with the 1F1B algorithm for concurrent forward and backward passes. A recent work, GNNPipe Chen et al. (2023) optimized GNN layer distribution across GPUs, reducing communication overhead and enabling hybrid parallelism for large-scale graphs.
3 Problem Formulation
@@ 3.1 Traffic Forecasting
The traffic forecasting task aims to predict the future traffic state $\mathcal{X}_{t+1:t+T^{\prime}}=[X_{t+1},\ldots,X_{t+T^{\prime}}]$ of $T^{\prime}$ future time steps from $t+1$ to $t+T^{\prime}$, based on historical traffic data $\mathcal{X}_{t-T+1:t}=[X_{t-T+1},\ldots,X_{t}]$ of $T$ time steps from $t-T+1$ to $t$. The objective is to learn a nonlinear mapping $f(\cdot)$, which captures the spatiotemporal dependencies from historical traffic observations to predict future traffic states:
$\mathcal{X}_{t-T+1:t}=[X_{t-T+1},\ldots,X_{t}]\xrightarrow{f(\cdot)}[X_{t+1},\ldots,X_{t+T^{\prime}}].$ (1)
3.2 Pipeline-Parallel Model Inference
Given a GNN model $\mathcal{M}$ for traffic forecasting, it consists of $L$ layers, denoted as $l_{j}$ (where $j=1,2,\dots,L$). Each layer $l_{j}$ is characterized by its memory consumption $\text{mem}_{j}$, input parameter size $\text{input}_{j}$, and output parameter size $\text{output}_{j}$. To facilitate a more streamlined inference process, we partition the model $\mathcal{M}$ into multiple stages and subdivide a mini-batch of size $B$ into multiple micro-batches, each of size $B_{\mu}$, allowing for better resource utilization and overlap of computation and communication. Each stage $\text{stage}_{k}$ is defined as:
$\text{stage}_{k}=\{l_{start_{k}},l_{end_{k}}\},\quad\bigcup_{k=1}^{K}\text{stage}_{k}=\mathcal{M}.$ (2)
@@ 3.2 Pipeline-Parallel Model Inference
$S$ =$=\{(d_{i},\text{stage}_{k})\mid i,k\in\{1,2,\dots,n\},\;n\leq N_{d};$ (3)
@@
Our goal is to find a joint strategy $(\mathcal{O},S)$, with the layer execution order $\mathcal{O}$ and the layer-device scheduling strategy $S$, that determines the partitioning of all layers of $\mathcal{M}$ into stages and their allocation to heterogeneous fog devices for conducting distributed parallel inference in a pipelined manner. $(\mathcal{O},S)$ leads to the two optimization objectives below.
3.2.1 Minimizing Longest Stage Execution Time
In pipeline parallelism, maximizing pipeline throughput is equivalent to minimizing the execution time of the longest stage. The execution time $T_{\text{exec}}(d_{i})$ for device $d_{i}$ involves computation time $T_{\text{comp}}(\text{stage}_{i},d_{i})$ and communication time $T_{\text{comm}}(\text{stage}_{i},d_{i-1},d_{i})$ for intermediate value transmission:
$T_{\text{exec}}(d_{i})=\max\left(\begin{aligned} &T_{\text{comp}}(\text{stage}_{i},d_{i}),\\ &T_{\text{comm}}(\text{stage}_{i},d_{i-1},d_{i})\end{aligned}\right),$ (4)
$T_{\text{comp}}(\text{stage}_{i},d_{i})=\sum_{l_{j}\in\text{stage}_{i}}\text{proc}_{i}(l_{j}),$ (5)
$T_{\text{comm}}(\text{stage}_{i},d_{i-1},d_{i})=\frac{\text{output}_{l_{end_{i}}}\cdot B_{\mu}}{\min(b_{i-1}^{\text{up}},b_{i}^{\text{down}})}.$ (6)
@@ 3.2.1 Minimizing Longest Stage Execution Time
3.2.2 Load Balancing
To mitigate pipeline inefficiencies caused by load imbalance, We aim to balance the load across devices by minimizing the standard deviation $\sigma$ of devices’ execution time $T_{\text{exec}}(d_{i})$:
$\sigma=\sqrt{\frac{1}{N_{d}}\sum_{i=1}^{N_{d}}\left(T_{\text{exec}}(d_{i})-\overline{T_{\text{exec}}}\right)^{2}}.$ (7)
3.2.3 Optimization Summary
In conclusion, to maximize the pipeline throughput, we optimize both the layer execution order $\mathcal{O}$ and the layer-device scheduling strategy $S$, searching for the optimal $(\mathcal{O}^{*},S^{*})$ that minimizes the maximum execution time across all devices:
$\min_{(\mathcal{O}^{*},S^{*})}\max_{1\leq i\leq N_{d}}\left\{\begin{aligned} \max\big{(}&T_{\text{comp}}(\text{stage}_{i},d_{i}),\\ &T_{\text{comm}}(\text{stage}_{i},d_{i-1},d_{i})\big{)}\end{aligned}\right\}.$ (8)
Figure 1: The network architecture of MultiGran-STGCN.
4 System Design: MultiGran-STGCNFog
This section begins with a system overview illustrating all phases of the proposed inference system. Next, MultiGran-STGCN components are elaborated in detail, along with the heterogeneous cross-device execution scheduling algorithm.
4.1 System overview
The workflow of MultiGran-STGCNFog consists of four phases: (1) Registration Phase: Fog nodes connect to the fog manager for node registration and initialization. (2) Profiling Phase: The fog manager distributes the forecasting model $\mathcal{M}$ and sampled historical traffic data to registered nodes, where inference is performed to profile computation time $\text{proc}_{i}(l_{j})$, memory usage $\text{mem}_{j}$, input and output parameter sizes, $\text{input}_{j}$ and $\text{output}_{j}$ for every layer, and P2P communication bandwidth $b_{i,i+1}$. (3) Scheduling Phase: The fog manager calculates the optimal scheme $(\mathcal{O}^{*},S^{*})$ and assigns stage allocation with corresponding model segments. (4) Runtime Phase: Fog nodes perform inference collaboratively, exchanging intermediate values as needed.
4.2 Traffic forecasting model
The architecture of MultiGran-STGCN is depicted in Fig. 1, and its four components will be introduced sequentially.
4.2.1 Multi-scale Traffic Data Modeling

### v2

# Forecasting at Full Spectrum: Holistic Multi-Granular Traffic Modeling under High-Throughput Inference Regimes

@@ Abstract.
Notably, current intelligent transportation systems rely heavily on accurate traffic forecasting and swift inference provision to make timely decisions. While Graph Convolutional Networks (GCNs) have shown benefits in modeling complex traffic dependencies, the existing GCN-based approaches cannot fully extract and fuse multi-granular spatiotemporal features across various spatial and temporal scales sufficiently in a complete manner, proven to yield less accurate results. Besides, as extracting multi-granular features across scales has been a promising strategy across domains such as computer vision, natural language processing, and time-series forecasting, pioneering studies have attempted to leverage a similar mechanism for spatiotemporal traffic data mining. However, additional feature extraction branches introduced in prior studies critically increased model complexity and extended inference time, making it challenging to provide fast forecasts. In this paper, we propose MultiGran-STGCNFog, an efficient fog distributed inference system with a novel traffic forecasting model that employs multi-granular spatiotemporal feature fusion on generated dynamic traffic graphs to fully capture interdependent traffic dynamics. The proposed scheduling algorithm GA-DPHDS, optimizing layer execution order and layer-device scheduling scheme simultaneously, contributes to considerable inference throughput improvement by coordinating heterogeneous fog devices in a pipelined manner. Extensive experiments on real-world datasets demonstrate the superiority of the proposed method over selected GCN baselines.
Traffic forecasting, Graph convolutional network, Multi-granular spatiotemporal feature fusion, Distributed inference system
copyright: noneconference: ; ; booktitle: doi: $c$s: Information systems Spatial-temporal systems
1. Introduction
@@
Nowadays, as a core pillar of smart city paradigm, intelligent transportation systems play a pivotal role in enhancing urban mobility, reducing congestion, and improving transportation efficiency. The large volume of traffic data has shifted attention toward data analysis algorithms and corresponding servings to enhance transportation system intelligence (Zantalis et al., 2019). Traditional centralized transportation systems are limited by high latency and weak scalability. Against this backdrop, fog computing, which processes data on nearer nodes close to the data source, has emerged as a more efficient and scalable solution. Collected traffic data needs to be analyzed efficiently to predict future traffic conditions as accurate predictions benefit services like transportation optimization, navigation planning, and congestion prevention (Li et al., 2022). Consequently, traffic forecasting has become an essential task, focusing on predicting future road network patterns using historical observations (Liu et al., 2023).
Embodied by advances across domains, the exploitation of multi-granular features at different scales has emerged as a promising mechanism for enhancing model expressiveness and robustness across a wide range of learning tasks, which demonstrates strong capability in capturing hierarchical structures and contextual dependencies from complex data. In computer vision, multi-scale designs are often realized through pyramid structures or parallel convolutional kernels, enabling models to detect objects and interpret scenes across spatial resolutions (Lin et al., 2017; Chen et al., 2018; Zhao et al., 2017; Szegedy et al., 2017). In natural language processing, hierarchical representations can be constructed by integrating local token-level dependencies with broader document-level semantics, via attention mechanisms that operate across multiple textual spans (Dai et al., 2019; Beltagy et al., 2020; Zhang et al., 2019; Jia et al., 2019; Jin et al., 2020). In the realm of time-series analysis, models benefit from capturing both short-term fluctuations and long-term temporal trends, typically through multi-resolution temporal blocks and frequency-domain decomposition (Lim et al., 2021; Zhou et al., 2021; Wu et al., 2022; Cui et al., 2016). In line with other research domains, the exploration of multi-scale mechanisms has commenced in traffic data mining.
@@
Existing traffic forecasting models primarily design complex Spatial-Temporal (S-T) blocks but lack a structured approach to extract and integrate hierarchical, multi-granular spatiotemporal features across spatial and temporal dimensions. Although a few works explored multi-granular spatial (Wang et al., 2021; Guo et al., 2021; Chen et al., 2024; Li et al., 2024) or temporal features (Dan et al., 2024; Wang and Jing, 2022; Huo et al., 2023; Guo et al., 2019) independently, traffic forecasting involves both spatial and temporal characteristics mining. Seamlessly integrating both is crucial for accurate forecasts and remains unexplored.
@@
Besides, previous studies manipulating multi-scale traffic data introduced multiple feature extraction branches to capture features of different granularities. This significantly increased model complexity and led to extended inference time, making it challenging for Graph Neural Network (GNN) serving where efficiency and low latency are critical. No solutions were proposed in these earlier works to address this issue and no framework was specifically designed for accelerating multi-granular spatiotemporal traffic forecasting.
While GCN-based models dominate spatiotemporal forecasting, transformer-based ones have recently shown efficacy in capturing long-term dependencies (Jiang et al., 2023; Liu et al., 2023). However, GCNs still remain irreplaceably unmatched in handling graph-structured spatial dependencies, and with carefully designed architectures, they are capable of achieving better performance than transformers (Dan et al., 2024; Zhang et al., 2024). In light of the above, we seek to answer two main research questions:
  1. RQ1.
    Whether constructing a comprehensive multi-scale feature extraction and fusion mechanism—jointly designed across both spatial and temporal dimensions—is able to enhance the accuracy of GCN-based traffic forecasting?
  2. RQ2.
    Can an inference-optimized framework benefit the inference throughput of multi-granular spatiotemporal traffic forecasting models, and in what manner does it achieve the improvement?
@@
Furthermore, designing accurate traffic forecasting models with swift inference still encounters several key challenges:
@@
1) Traffic data exhibits highly dynamic and sophisticated patterns, making it difficult to model spatiotemporal correlations.
@@
2) In GNN serving systems, efficiently leveraging available resources is essential for speeding up where communication latency becomes a major hurdle (Zeng et al., 2022).
@@
3) Optimizing distributed parallel inference schemes is increasingly complicated due to the exponentially-growing search space for the optimal layer execution order and layer allocation scheduling, together with layer input-output dependencies.
@@
4) Mainstream distributed parallel frameworks assume even partitioning of stages, ignoring hardware heterogeneity. Real-word device clusters are often heterogeneous, and model layers cannot be fully in even distribution to ensure equal processing times.
@@
To fill these gaps, we propose MultiGran-STGCNFog, an efficient GNN inference system that extracts and fuses multi-granular spatiotemporal features across various spatial and temporal scales. The distributed pipeline-parallel architecture across fog nodes enables MultiGran-STGCNFog to tackle the extended inference latency, supporting accurate and swift traffic forecasting.
@@
  • A novel GCN, Multi-Granularity Spatiotemporal Graph Convolution Network (MultiGran-STGCN), is proposed. It employs dynamic graph techniques and a multi-granular spatiotemporal feature fusion mechanism through Laplacian-driven hierarchical graph clustering and multi-scale temporal modeling to exploit multi-granular features.
@@
  • To mitigate the prolonged inference time induced by additional feature extraction branches, we construct MultiGran-STGCNFog for inference acceleration, leveraging the resources of heterogeneous fog devices. It enables faster inference by pipeline parallelism, which allows immediate data transmission among fog nodes in proximity to avoid remote transfers.
@@
  • We design a robust heterogeneous cross-device scheduling algorithm, GA-DPHDS, to optimize the layer execution order and layer-device scheduling scheme jointly, overcoming the search space explosion challenge. Experiment results show the necessity of layer execution order optimization, which is overlooked in previous pipeline parallelism works.
@@
  • The proposed framework is tested with physical computing nodes and evaluated using three real-world traffic datasets. Up to 9.86% forecasting metric improvement and 2.43x throughput improvement demonstrate its superior performance compared with selected baselines.
2. Related Work
2.1. GCNs: Spatiotemporal Traffic Forecasting
@@
When combining temporal modeling techniques like RNNs, GCN-based models are particularly effective at capturing spatiotemporal dependencies (Seo et al., 2018; Cui et al., 2020). DCRNN (Li et al., 2017) modeled traffic as a diffusion process on directed graphs, using bidirectional diffusion convolution and RNNs to outperform baselines. Gated-STGCN (Yu et al., 2017) leveraged graph convolutions for spatial and gated 1D convolutions for temporal modeling, improving efficiency by removing recurrent structures. ASTGCN (Guo et al., 2019) employed attention mechanisms to capture spatial and temporal dependencies, while HGCN (Guo et al., 2021) modeled traffic across micro- and macro-level networks. GWNET (Wu et al., 2019) combined dilated causal and graph convolutions for hidden dependency capture, achieving state-of-the-art results, and OGCRNN (Guo et al., 2020) integrated GCN with recurrent units for enhanced feature extraction.
2.2. Distributed and Parallel GNNs
@@
The growing computational demands of deep learning and the inability to deliver fast serving drove a shift from centralized to distributed computing to leverage parallelism in GNNs. Data parallelism partitions input data across nodes, using graph parallelism to minimize inter-node communication or mini-batch parallelism with node sampling techniques. Model parallelism distributes the model itself through operator parallelism with simultaneous operations, pipeline parallelism which processes stages sequentially, or ANN parallelism to spread layers across devices (Besta and Hoefler, 2024).
@@
GPipe (Huang et al., 2019) applied pipeline parallelism by dividing models into stages across accelerators for micro-batch processing, while PipeDream (Narayanan et al., 2019) enhanced this with the 1F1B algorithm for concurrent forward and backward passes. A recent work, GNNPipe (Chen et al., 2023) optimized GNN layer distribution across GPUs, reducing communication overhead and enabling hybrid parallelism for large scales.
3. Problem Formulation
@@ 3.1. Traffic Forecasting
The traffic forecasting task aims to predict the future traffic state $\mathcal{X}_{t+1:t+T^{\prime}}=[X_{t+1},\ldots,X_{t+T^{\prime}}]$ of $T^{\prime}$ future time steps from $t+1$ to $t+T^{\prime}$, based on historical traffic data $\mathcal{X}_{t-T+1:t}=[X_{t-T+1},\ldots,X_{t}]$ of $T$ time steps from tT+1t-T+1 to $t$. The objective is to learn a nonlinear mapping $f(\cdot)$, which captures the spatiotemporal dependencies from historical traffic observations to predict future traffic states:
(1) $\mathcal{X}_{t-T+1:t}=[X_{t-T+1},\ldots,X_{t}]\xrightarrow{f(\cdot)}[X_{t+1},\ldots,X_{t+T^{\prime}}].$
3.2. Pipeline-Parallel Model Inference
Given a GNN model $\mathcal{M}$ that consists of $L$ layers for traffic forecasting, denoted as $l_{j}$ (where $j=1,2,\dots,L$). Each layer $l_{j}$ is characterized by its memory consumption $\text{mem}_{j}$, input parameter size $\text{input}_{j}$, and output parameter size $\text{output}_{j}$. To facilitate a more streamlined inference process, we partition $\mathcal{M}$ into multiple stages and subdivide a mini-batch of size $B$ into multiple micro-batches, each of size $B_{\mu}$, allowing for better resource utilization and overlap of computation and communication. Each $\text{stage}_{k}$ is defined as:
(2) $\text{stage}_{k}=\{l_{start_{k}},l_{end_{k}}\},\quad\bigcup_{k=1}^{K}\text{stage}_{k}=\mathcal{M}.$
@@ 3.2. Pipeline-Parallel Model Inference
(3) $S$ $=\{(d_{i},\text{stage}_{k})\mid i,k\in\{1,2,\dots,n\},\;n\leq N_{d};$
@@
Our goal is to find a joint strategy $(\mathcal{O},S)$, with layer execution order $\mathcal{O}$ and layer-device scheduling strategy $S$, that determines the partitioning of all layers of $\mathcal{M}$ into stages and their allocation to heterogeneous fog devices for distributed parallel inference in a pipelined manner. $(\mathcal{O},S)$ leads to two optimization objectives.
3.2.1. Minimizing Longest Stage Execution Time
In pipeline parallelism, maximizing pipeline throughput is equivalent to minimizing the execution time of the longest stage. The execution time $T_{\text{exec}}(d_{i})$ for device $d_{i}$ involves computation time $T_{\text{comp}}(\text{stage}_{i},d_{i})$ and intermediate value transmission time $T_{\text{comm}}(\text{stage}_{i},d_{i-1},d_{i})$:
(4) $T_{\text{exec}}(d_{i})=\max\left(\begin{aligned} &T_{\text{comp}}(\text{stage}_{i},d_{i}),\\ &T_{\text{comm}}(\text{stage}_{i},d_{i-1},d_{i})\end{aligned}\right),$
(5) $T_{\text{comp}}(\text{stage}_{i},d_{i})=\sum_{l_{j}\in\text{stage}_{i}}\text{proc}_{i}(l_{j}),$
(6) $T_{\text{comm}}(\text{stage}_{i},d_{i-1},d_{i})=\frac{\text{output}_{l_{end_{i}}}\cdot B_{\mu}}{\min(b_{i-1}^{\text{up}},b_{i}^{\text{down}})}.$
@@ 3.2.1. Minimizing Longest Stage Execution Time
3.2.2. Load Balancing
To mitigate pipeline inefficiencies caused by load imbalance, which introduces idle times on faster devices, we aim to balance the load across devices by minimizing the standard deviation $\sigma$ of devices’ execution time $T_{\text{exec}}(d_{i})$:
(7) $\sigma=\sqrt{\frac{1}{N_{d}}\sum_{i=1}^{N_{d}}\left(T_{\text{exec}}(d_{i})-\overline{T_{\text{exec}}}\right)^{2}}.$
3.2.3. Optimization Summary
In conclusion, to maximize the pipeline throughput, we optimize both the layer execution order $\mathcal{O}$ and layer-device scheduling $S$, searching for the optimal $(\mathcal{O}^{*},S^{*})$ that minimizes the maximum execution time across all devices:
(8) $\min_{(\mathcal{O}^{*},S^{*})}\max_{1\leq i\leq N_{d}}\left\{\begin{aligned} \max\big{(}&T_{\text{comp}}(\text{stage}_{i},d_{i}),\\ &T_{\text{comm}}(\text{stage}_{i},d_{i-1},d_{i})\big{)}\end{aligned}\right\}.$
Figure 1. The overview of MultiGran-STGCNFog.
[…]
System overall diagram.
Figure 2. The network architecture of MultiGran-STGCN.
[…]
Network architecture.
4. System Design: MultiGran-STGCNFog
This section begins with a system overview illustrating all phases of the proposed inference system. Next, MultiGran-STGCN components are elaborated in detail, along with the heterogeneous cross-device execution scheduling algorithm. Lastly, a theoretical analysis establishing a lower bound on throughput improvement across varying datasets is conducted to demonstrate the robustness.
4.1. System overview
The workflow of MultiGran-STGCNFog consists of four phases shown in Fig. 1: (1) Registration Phase: Fog nodes connect to the fog manager for node registration and initialization. (2) Profiling Phase: The fog manager distributes the forecasting model $\mathcal{M}$ and sampled historical traffic data to registered nodes, where inference is performed to profile computation time $\text{proc}_{i}(l_{j})$, memory usage $\text{mem}_{j}$, input and output parameter sizes, $\text{input}_{j}$ & $\text{output}_{j}$ for every layer, and P2P communication bandwidth $b_{i,i+1}$. (3) Scheduling Phase: The fog manager calculates the optimal scheme $(\mathcal{O}^{*},S^{*})$ and assigns stage allocation with corresponding model segments. (4) Runtime Phase: Fog nodes perform inference collaboratively, exchanging intermediate values as needed.
4.2. Traffic forecasting model
The architecture of MultiGran-STGCN is depicted in Fig. 2, and its four components will be introduced sequentially in this section.
4.2.1. Multi-scale Traffic Data Modeling